a stone thrown down with a speed U take a time T1 to reach the ground while another stone thrown upward from the same point with same speed takes time T2.the maximum height the II stone reaches from the ground is

Grade:12

1 Answers

Tashyab
5 years ago
For the given question,Let the dist. travelled by the stone thrown down be h.So, h = uT1 + 1/2 g(T1)^2Now, for the stone thrown upwards,Initial velocity= -u and a=gAnd let the dist. travelled by it to the max point from the point of throwing be H. So that total dist. covered by the 2nd thrown is H+h.So, H+h= -uT2 +1/2g(T2)^2Now, H+h - h = [-uT2 + 1/2 g(T2)^2 ] - [uT1 + 1/2g(T1)^2]=> H = -uT2 - uT1 + 1/2g(T2)^2 - 1/2g(T1)^2=> H = -u ( T2 - T1 ) + 1/2g ( (T2)^2 - (T1)^2 ) OR=> H = u ( T1 - T2 ) - 1/2g ( (T1)^2 - (T2)^2 )Thank you!!